OpenStudy (anonymous):

Evaluate cot 45° without using a calculator by using ratios in a reference triangle.

OpenStudy (blurbendy):

cot(45) = cos(45) / sin(45) cos(45) = sin (45) So, cot(45) is equal to 1

OpenStudy (blurbendy):

cos (45) = sqrt[2] / 2 sin (45) = sqrt[2] / 2

OpenStudy (anonymous):

Theres more? I thought that was it D:

OpenStudy (blurbendy):

no, im just showing you why cos(45) = sin(45)

OpenStudy (anonymous):

Oh, Okay! Well thank you!

OpenStudy (blurbendy):

yw